# Prompting LLMs for Code Editing: Struggles and Remedies

### v1

@@ 7. Threats to Validity
The use of one month of data may be insufficient for broad generalization, particularly given the rapid evolution of AI tools at Google, and the corresponding adaptation of developers. Furthermore, updates to the LLMs supporting the AI tools, which occur during tool evolution, might influence developer satisfaction and behavior. However, we had to tolerate these limitations to balance data collection complexity, privacy considerations, and the scale of analysis. Given the substantial size of Google’s internal developer base and the widespread adoption of Transform Code among developers over an extended period, we believe that our dataset is representative, but should be interpreted considering the use case, user population, and the scope of features.
@@
The error analysis is subject to the typical threats to validity encountered in other qualitative analyses. While analyzing data qualitatively allows us to engage in-depth and distill knowledge from previously unexplored areas, we cannot guarantee that the findings generalize beyond the sampled examples without further validation. Specifically, because complex and lengthy examples were excluded to facilitate practical manual review, our sample may not cover all possible prompts that are used for code editing with an AI tool. However, given that Google covers diverse application domains and developers with varying expertise in programming, domain knowledge, and AI tools, we believe the sampled examples reasonably represent diverse use cases of code editing tasks.
@@
Our experiments with improving prompt effectiveness have all been in the context of Transform Code. We believe that the ideas will carry over to similar tools where prompts are used to make code edits (e.g. Cursor (cur, 2025)).
@@ 8. Discussion & Conclusion
Our approach is to analyze the common reasons—the categories of missing detail—that the prompts themselves could be lacking. We did this by leveraging an LLM and asking it to articulate the prompt enhancements that would be needed to get Transform Code to produce edits that match the given desired edit that each of those prompts was supposed to produce. We collated this information into a codebook that captures five common classes of missing information: missing some specific information, missing an operationalization plan, faulty localization/scope, missing codebase context, or missing intent.
This analysis then enabled us to create AutoPrompter, a tool that pro-actively adds details to the prompt (without knowing the ground truth, of course) to get Transform Code to produce a more satisfactory edit from a given developer prompt. This idea seems to have shown promise, as we show in our test set: 27% of the cases that were not achieving a satisfactory result with Transform Code now accomplished a satisfactory result with the automatically enhanced prompt.
@@
Another promising usage of AutoPrompter would be to use it for the automatic curation of training data from the usage logs. Transform Code usage data can be very useful for training, as it offers negative examples (original Transform Code edit) alongside ground-truths (user ”rewrites”) given user prompts. However, only a small, human-verified subset of user-submitted ground truths (see Section 4.1) has been used, due to the challenges in extracting quality ground-truths, discussed in Section 4.2. The core idea is to include the AutoPrompter-improved Transform Code edit (edit generated using the AutoPrompter-enhanced user prompt) as ground-truths, only if they are identical to a future snapshot. Because AutoPrompter only augments the prompt based on the provided context, the improved Transform Code edit captures the original user intent without adding requirements not mentioned in the original prompt. The identity to a future code snapshot also shows implicit user confirmation that the improved edit was desired.
Finally, one potential benefit of AutoPrompter, not discussed in this paper as it is out of scope, is the increased transparency and control it offers users. As an enhanced prompt is generated to obtain an enhanced code edit from Transform Code, when presented, users can easily verify whether their intent is accurately described in the prompt. This prompt can then be easily updated by the user if it contains errors, affording them direct control over the input to Transform Code. In addition, presenting these improved prompts to developers could help them learn to write more effective prompts themselves (Ma et al., 2024), possibly enabling them to understand the benefits of adding more context. This understanding, in turn, could lead to better code edits from Transform Code with less accumulated effort.

### v2

@@ 7. Threats to Validity
The use of one month of data may be insufficient for broad generalization, particularly given the rapid evolution of AI tools at Google, and the corresponding adaptation of developers. Furthermore, updates to the LLMs supporting the AI tools, which occur during tool evolution, might influence developer satisfaction and behavior. However, we had to tolerate these limitations to balance the complexity of data collection, privacy considerations, and the scale of analysis. Given the substantial size of Google’s internal developer base and the widespread adoption of Transform Code among developers over an extended period, we believe that our dataset is representative, but should be interpreted considering the use case, user population, and the scope of features.
@@
The error analysis is subject to the typical threats to validity encountered in other qualitative analyses. While analyzing data qualitatively allows us to engage in-depth and distill knowledge from previously unexplored areas, we cannot guarantee that the findings generalize beyond the sampled examples without further validation. Specifically, because complex and lengthy examples were excluded to facilitate practical manual review, our sample may not encompass all possible prompts used for code editing with an AI tool. However, given that Google covers diverse application domains and developers with varying expertise in programming, domain knowledge, and AI tools, we believe the sampled examples reasonably represent a diverse range of use cases for code editing tasks.
@@
Our experiments with improving prompt effectiveness have all been in the context of Transform Code. We believe that the ideas will carry over to similar tools where prompts are used to make code edits (e.g., Cursor (cursor)).
@@ 8. Discussion & Conclusion
Our approach is to analyze the common reasons that the prompts themselves could be lacking. We did this by leveraging an LLM and asking it to articulate the prompt enhancements that would be needed to get Transform Code to produce edits matching the given desired edit. We collated this information into a codebook that captures five common classes of missing information: missing some specific information, missing an operationalization plan, faulty localization/scope, missing codebase context, or missing intent.
This analysis then enabled us to create AutoPrompter, a tool that proactively adds details to the prompt, without knowing the ground truth, to get Transform Code to produce a more satisfactory edit from a given developer prompt. This idea seems to have shown promise, as we show in our test set: 27% of the cases that were not achieving a satisfactory result with Transform Code now accomplished a satisfactory result with the automatically enhanced prompt.
@@
Another promising usage of AutoPrompter would be to use it for the automatic curation of training data from the usage logs. Transform Code usage data can be very useful for training, as it offers negative examples (original Transform Code edit) alongside ground-truths (user “rewrites”) given user prompts. However, only a small, human-verified subset of user-submitted ground truths (see Section 4.1) has been used, due to the challenges in extracting quality ground-truths, discussed in Section 4.2. The core idea is to include the AutoPrompter-improved Transform Code edit (edit generated using the AutoPrompter-enhanced user prompt) as ground-truths, only if they are identical to a future snapshot. Because AutoPrompter only augments the prompt based on the provided context, the improved Transform Code edit captures the original user intent without adding requirements not mentioned in the original prompt. The identity of a future code snapshot also shows implicit user confirmation that the improved edit was desired.
Finally, one potential benefit of AutoPrompter, not discussed in this paper as it is out of scope, is the increased transparency and control it offers users. As an enhanced prompt is generated to obtain an enhanced code edit from Transform Code, when presented, users can easily verify whether their intent is accurately described in the prompt. This prompt can then be easily updated by the user if it contains errors, affording them direct control over the input to Transform Code. In addition, presenting these improved prompts to developers could help them learn to write more effective prompts themselves (ma2024engineer), possibly enabling them to understand the benefits of adding more context. This understanding, in turn, could lead to better code edits from Transform Code with less accumulated effort.